Out of 100 students, (49 girls, 51 boys) if three children are selected what is the probability that at least one is a girl?
I am missing a key point to this equation because I keep getting 49/97 which I believe is wrong
Thanks for any help

Found 2 solutions by stanbon, oberobic:
Answer by stanbon(75887)

Out of 100 students, (49 girls, 51 boys) if three children are selected what is the probability that at least one is a girl?
--------
P(at least one girl) = 1 - P(3 boys)
---
= 1 - 51C3/100C3
---
= 1 - 0.1288
---
= 0.8712
==============
Cheers,
Stan H.
==============

Answer by oberobic(2304)

The probability of picking a girl on the first pick is: 49/100.
.
The probability of picking a girl on the second pick depends on whether a girl was picked first.
If a girl was picked initially, then the probability of picking a girl on the second pick is : 48/99.
If a boy was picked initially, then the probability of picking a girl on the second pick is: 49/99.
.
The probability of picking a girl first and second is: 49/100 * 48/99.
.
And so forth.
.
You need to determine the conditional probabilities to find the chance of picking at least 1 girl.
.
1st Pick      2nd Pick      3rd Pick         Probability             Description          
P(girl)=49/100      P(girl)=48/99      P(girl)=47/98      P=0.113939394 (A) 3 girls
" " P(boy)=51/98 P=0.123636364 (B) 2 girls, 1 boy
" P(boy)=51/99 P(girl)=48/98 P=0.123636364 (C) 2 girls, 1 boy
" " P(boy)=50/98 P=0.128787879 (D) 1 girl, 2 boys
P(boy)=51/100 P(girl)=49/99 P(girl)=48/98 P=0.123636364 (E) 2 girls, 1 boy
" " P(boy)=50/98 P=0.128787879 (F) 1 girl, 2 boys
" P(boy)=50/99 P(girl)=49/98 P=0.128787879 (G) 1 girl, 2 boys
" " P(boy)=49/98 P=0.128787879 (H) 3 boys

The total probability = 1.00, as you can see.
.
P(at least 1 girl) = A+B+C+D+E+F+G = 0.871212121
.
OR
.
You could see the problem in terms of the chance of picking 3 boys. If you do not pick 3 boys, then you have to pick at least 1 girl.
.
P(picking 3 boys) = 51/100 * 50/99 * 49/98 = 0.128787879
.
P(picking at least 1 girl) = 1.0 - 0.128787879 = 0.871212121
.
Same answer. Two ways to get there.
.
Done.